A 25% chlorine solution was mixed with an 18% solution to create a 20% solution. How much of the 18% and 25% solution were used to make 28L of the 20% solution?
Answer by stanbon(75887)

A 25% chlorine solution was mixed with an 18% solution to create a 20% solution. How much of the 18% and 25% solution were used to make 28L of the 20% solution?
----------------------
Equation:
active ingredient + active ingre = active ingre
0.25x + 0.18(28-x) = 0.20*28
Multiply thru by 100 to get:
25x + 18*28 -18x = 20*28
7x = 2*28
x = 8L (amt. of 25% solution in the mixture)
28-x = 28-8 = 20 L (amt. of 18% solution in the mixture)
===========================================================
Cheers,
Stan H.